### v1

# Clue-Guided Path Exploration: An Efficient Knowledge Base Question-Answering Framework with Low Computational Resource Consumption

Abstract
In recent times, large language models (LLMs) have showcased remarkable capabilities. However, updating their knowledge poses challenges, potentially leading to inaccuracies when confronted with unfamiliar queries. While integrating knowledge graphs with LLMs has been explored, existing approaches treat LLMs as primary decision-makers, imposing high demands on their capabilities. This is particularly unsuitable for LLMs with lower computational costs and relatively poorer performance. In this paper, we introduce a Clue-Guided Path Exploration framework (CGPE) that efficiently merges a knowledge base with an LLM, placing less stringent requirements on the model’s capabilities. Inspired by the method humans use to manually retrieve knowledge, CGPE employs information from the question as clues to systematically explore the required knowledge path within the knowledge base. Experiments on open-source datasets reveal that CGPE outperforms previous methods and is highly applicable to LLMs with fewer parameters. In some instances, even ChatGLM3, with its 6 billion parameters, can rival the performance of GPT-4. Furthermore, the results indicate a minimal invocation frequency of CGPE on LLMs, suggesting reduced computational overhead. For organizations and individuals facing constraints in computational resources, our research offers significant practical value.
1 Introduction
Large language models (LLMs) Achiam et al. (2023); Ouyang et al. (2022); Touvron et al. (2023); Cheng et al. (2023) have shown their power on many natural language process tasks. Nevertheless, there are certain issues associated with LLMs when undertaking question-answering tasks. The knowledge of LLMs remains static, frozen at the point of their training, making it challenging to incorporate timely updates. Furthermore, when faced with unfamiliar queries, LLMs may sometimes provide incorrect answers with an overconfident tone Luo et al. (2023), a phenomenon often referred to as hallucination Bang et al. (2023); Ji et al. (2023). Acknowledging these challenges, an inherent and auspicious resolution involves the integration of external knowledge Mialon et al. (2023), specifically employing knowledge graphs (KGs), to enhance the question-answering proficiency of LLMs.
Figure 1: Comparison among three different types of methods. The dashed box indicates the capability requirements of the methods for LLMs. The first two types of methods have higher demands on the overall capabilities of LLMs, requiring them to complete comprehensive tasks. In contrast, our proposed clue-guided path exploration method has relatively lower requirements on the overall capabilities of LLMs, only demanding them to complete a one task at a time.
Researchers have conducted some studies to combine external knowledge and LLMs. Some researchers prompt LLMs to generate database operation instructions to retrieve information Hu et al. (2023); Wang et al. (2023). Chain-of-thought Wei et al. (2022b) is also used to help LLMs decompose the input question into multiple sub-questions Li et al. (2023a); Zhao et al. (2023). However, in these methods, LLMs lack awareness of the nature of external knowledge sources. As a result, the instructions or sub-questions generated by LLMs may not align well with the external knowledge. TOG Sun et al. (2023b) addresses the mismatch issue between LLMs and knowledge bases by facilitating multi-turn communications, continually supplying updated knowledge path information to LLMs. Nevertheless, in the course of TOG’s path exploration, critical decisions, such as choosing relationships, selecting entities, and determining termination points, heavily depend on the inherent capabilities of the large language model. TOG lacks the inclusion of any mechanism or provision of supplementary information to aid LLMs in decision-making. This presents a significant challenge to the overall capabilities of LLMs and impedes the accurate exploration of knowledge paths. As a result, it makes the approach less suitable for LLMs with comparatively weaker comprehensive abilities. This issue persists and remains a common challenge faced by previous methods. As the capabilities of LLMs strengthen, their parameter sizes increase, necessitating more computational resources, such as GPUs. For a substantial proportion of organizations and individuals, the availability of computational resources is a crucial consideration. Therefore, developing a mechanism to lower the demands on LLMs’ capabilities during path exploration not only enhances the effectiveness of knowledge base question-answering but also carries significant practical value.
In everyday scenarios, when confronted with the task of retrieving information from a knowledge base, such as a book, in response to a question, we typically initiate the search process by distilling pertinent details from the inquiry. These extracted fragments serve as clues, enabling us to methodically narrow down the search scope and ultimately arrive at the desired answer. Each step in this process is straightforward yet remarkably effective. Drawing inspiration from this practical approach, we have developed a Clue-Guided Path Exploration (CGPE) framework. This framework harnesses the information embedded in questions as clues, guiding a systematic exploration of the knowledge graph and progressively revealing the path to the requisite knowledge.
For a given question, our methodology initially extracts ‘clues’ from the question. Based on these clues, we identify initial nodes of knowledge paths within the knowledge base. With the initial nodes, we can retrieve the related nodes. Then we match the information of these related nodes with clues to determine the next node for each path. We iterate through this process until all clues have undergone a matching cycle. We iterate through this process until we find a path that matches all the clues. This refined knowledge path is then presented to LLMs for answering the question. Notably, our approach positions the clues as guiding indicators throughout the entire path exploration process. In doing so, LLMs only need to focus on information matching using the given clues and information, without the added complexity of analyzing questions at the same time. This minimizes the requirement for LLMs to possess extensive capabilities, thereby improving the efficiency of path exploration and making our approach suitable for LLMs with fewer parameters.
The main contributions of this paper are summarized as follows:
  • We present CGPE, a novel clue-guided path exploration framework aimed at assisting large language models (LLMs) in answering questions. By incorporating a path exploration mechanism centered around clues, we have effectively lower the demands on LLMs’ capabilities.
  • Experiments conducted on various datasets demonstrate that our proposed method outperforms the previous state-of-the-art method. Notably, our approach performs well on LLMs with fewer parameters, marking a significant breakthrough compared to previous methods.
  • Experiments show that our method involves very few invocations of LLMs, significantly reducing computational resource consumption. Ablation studies offer insights for further reducing the consumption of computational resources.
2 Related Work
2.1 Knowledge Enhanced Large Language Models
One direct approach to addressing the generation of illusions by LLMs is the use of external knowledgeMialon et al. (2023) to enhance the model’s ability to perceive the real world. This has led to the emergence of numerous knowledge-based models. Information retrieval is commonly employed in developing knowledge-based chatbotsShuster et al. (2021). BlenderBot2Xu et al. (2021) integrates Internet search capabilities. SeeKeRShuster et al. (2022a) surpasses BlenderBot2Xu et al. (2021) by executing three modular tasks using a single language model: generating search queries, extracting relevant knowledge from retrieved documents, and generating the final response. BlenderBot3Shuster et al. (2022b) fine-tunes a 175-billion-parameter OPTZhang et al. (2022) on a combination of 20 question-answering and dialogue datasets. AtlasIzacard et al. (2023) stands out as a top model on the KILT benchmarkPetroni et al. (2021), which includes 11 knowledge-oriented tasks, including Wizard of WikipediaDinan et al. (2019).
2.2 Knowledge base Enhanced Models
There are diverse ways to enhance models based on knowledge, aiming to address issues arising from complex reasoning, hallucination, and poor interpretability. Cao et alCao et al. (2021) improved implicit event causality extraction by leveraging structured knowledge bases. Zhao et alZhao et al. (2023) recently proposed a Verification and Editing (VE) framework. This framework improves the Chain of Thought (CoT) reasoning in large language models (LLMs) by integrating retrieval systemsWei et al. (2022a). Researchers have also introduced the Chain-of-Knowledge (CoK)Li et al. (2023b), a novel framework designed to enhance the factual correctness of LLMs using heterogeneous knowledge sources. This ensures that inaccuracies in the foundational principles do not propagate to subsequent steps. Think-on-Graph(ToG) Sun et al. (2023a) is a plug-and-play framework that seamlessly applies to various LLMs and KGs. ToG also enhances the reasoning capabilities, enabling smaller LLMs (such as LLAMA2-70B) to compete with larger LLMs (such as GPT-4).
3 Method
As shown in Figure 2, our framework can be divided into three stages: clue extraction stage, path exploration stage, and answering stage. The clue extraction stage is elaborated in Section 3.2, while the path exploration stage is discussed in Sections 3.3, 3.4, and 3.5. As for the answering stage, we did not implement any specific design but rather provided the extracted knowledge to Large Language Models in the form of triplets to assist them in answering questions.
Figure 2: The workflow of our proposed CGPE framework. The green box indicates that these steps require the use of LLMs. The process of exploring the next node is within the dashed box, and it continues until the termination condition is met.
3.1 Problem Formulation
For a given question, we initially extract a set of clues, denoted as $C=\{c_{1},c_{2},\ldots,c_{n}\}$. We employed $P=\{p_{1},p_{2},\ldots,p_{m}\}$ to denote all paths explored within the knowledge base, where each $p_{i}$ represents an individual path encompassing all entities along that path. Each node potentially corresponds to a clue $c_{i}$ from set $C$. The clues corresponding to the nodes are marked as ‘used’ for the path, and they will not be used again to explore this path. Our objective is to identify a path that corresponds to all clues $c_{i}$ within set $C$. Ultimately, we present this curated path to the LLM to facilitate the answering of the given question.
3.2 Extracting Clues from the Given Question
Intuitively, we posit that the majority of information (not all information) within a question can be characterized by entities and relationships originating from the question. Particularly, the knowledge required to answer a question is presumed to be related with these entities and relationships. Employing these entities and relationships as clues for knowledge retrieval aids in addressing the given question. For instance, in the case of the question “What is the population of the capital of China?” utilizing (China, capital, population) as clues facilitates the effortless retrieval of relevant knowledge to answer the question.
The delineation between entities and relationships is not always clear-cut. In the aforementioned example, ‘capital’ can function both as a relationship and an entity. Additionally, relationships can, to some extent, be represented and inferred by entities Bordes et al. (2013); Sun et al. (2019). For instance, with an incomplete triple (China, ?, Beijing), one can infer the missing relationship as ‘capital’. Consequently, in our practical extraction approach, we do not rigidly distinguish between entities and relationships. Instead, we broadly task the LLM with extracting all entities from the question, utilizing them as clues to guide the exploration of paths within the knowledge graph. In this scenario, words that can function both as entities and relationships are also extracted.
3.3 Selection of Initial Nodes
Upon extracting clues from a given question, we initiate a path exploration within the knowledge graph. To choose the initial nodes for the paths, we conduct node queries in the knowledge graph using clues within the set $C$ as the value of nodes. In the scenario where the knowledge graph contains the necessary information to answer questions, we typically achieve successful queries for a specific clue (i.e. node). there invariably exists certain clue $C_{i}$ that can be retrieved, corresponding to specific nodes. These nodes serve as the initial nodes for paths, with the number of paths aligning with the quantity of nodes. And the corresponding clues are marked as ‘used’ for each path.
It is important to note that although these initial nodes mark the inception of a path, they may not necessarily be the commencement of the target knowledge chain; they could also occupy an intermediary position within the desired knowledge chain. This may pose a challenge for path exploration. We will address this issue in subsequent steps.
3.4 Exploration of the Next Node
For each path, we iteratively explore the next node until the nodes along the path collectively cover all clues in set $C$. Employing a predefined query for retrieving related nodes of the current node, we obtain a candidate set of these related nodes, capturing relationships, entities, and entity labels. The labels serve as abstract summaries of entity information, aligning with the abstract and non-specific nature in which clues are often presented in questions. Most knowledge graphs label entities, and if a label is absent, it is treated as ‘none’. Subsequently, we proceed through a four-step process to match the candidate set with clues.
Character Matching. Just as discussed when selecting the initial node, the initial node for a path is matched with the characters of a certain clue. If a related node matches an unused clue, then we consider this node as the next one in the path. And the clue $c_{i}$ is marked as ‘used’ for the path.
Fuzzy Matching. If no node is selected after character matching, we then prompt the LLM to conduct a fuzzing matching. Nodes in the candidate set are characterized with (relationship, label), which is a highly abstract representation of node information. We prompt LLM to match (relationship, label) with unused clues. If the model deems (relationship, label) as unrelated to any unused clue, the corresponding entities are pruned from the candidate set. Note that different nodes may be characterized with the same (relationship, label).
Rough Matching. The remaining nodes in the candidate set are then represented as (label, entity), in which the label is used to provide abstract information and the entity is used to provide specific information about the node. We prompt LLM to match and score the (label, entity) with unused clues. Considering that non-fine-tuned LLMs may not provide highly accurate scores, we simply remove nodes with a score of 0. But with a fine-tuned LLM, selecting the top-k nodes can reduce the frequency of invoking the LLM during the exact matching step.
Exact Matching. We perform exact matching on the remaining nodes to ultimately determine the next node in the path. The remaining nodes are still represented as (label, entity), and we prompt LLM to match each (label, entity) with unused clues. For each (label, entity), the most compatible clue $c_{i}$ is marked as ‘used’, and the corresponding node is chosen as the next node. If the LLM does not find a matching clue for the (label, entity), then the corresponding node is discarded. If multiple nodes are chosen as the next node, this step generates multiple new paths. Considering the scenario where initial nodes may lie within the middle of the target knowledge chain, theoretically, two paths exploring both ends of the target knowledge chain may be generated.
The rough matching and exact matching steps share similarities in terms of the information provided to the LLM. The distinction lies in their respective emphases, with the former prioritizing screening and the latter emphasizing precision. This division is deemed advantageous for achieving favorable outcomes on relatively less capable LLMs. Given the relatively low computational cost of these LLMs, the trade-off between multiple computations and enhanced effectiveness is deemed worthwhile. For LLMs with high capabilities and significant computational demands, keeping only the rough matching step is also acceptable.
3.5 Termination of Exploration
If a path corresponds to all clues $c_{i}$ within set $C$, it is considered to encompass all the necessary knowledge to answer the question. We then terminate the exploration of this path and use it as a knowledge chain to help the LLM answer the given question. In instances where paths still possess unused clues, yet fail to identify the next nodes, exploration for those paths is also terminated. Such paths can be categorized into two distinct types. The first is an incorrect path that ought to be disregarded. The second entails the correct path, wherein the initial node of the path is situated at the midpoint of the target knowledge chain, thereby leading to the exploration of nodes in a unidirectional manner. As previously discussed, another path theoretically exists, exploring the opposite direction. Merging these two paths usually ensures coverage of all clues in set $C$.
Upon the discovery of a path containing the requisite knowledge to answer the question, exploration of other paths is halted. The identified path is then presented to the LLM in the form of triplets, helping the LLM to answer the given question.
4 Experiments
4.1 Experimental Settings
[…]
Baselines. We compared our approach with three baseline methods: standard prompting (IO prompt) Brown et al. (2020), chain of thought prompting (CoT prompt) Wei et al. (2022b) and TOG Sun et al. (2023b). The knowledge base and Q&A sets in MOOC Q&A are highly interconnected, with most questions requiring knowledge retrieval from the knowledge base for answers. IO prompt and CoT prompt are two methods without external knowledge, and they are used to illustrate the proportion of questions requiring knowledge. TOG is a previous state-of-the-art method in knowledge base question-answering, used to validate the effectiveness of our knowledge retrieval approach.
Experiment Details. We conducted experiments on three different LLMs, namely ChatGLM3 Du et al. (2022) with 6 billion parameters, ChatGPT, and GPT-4. ChatGPT and GPT-4 are called by OpenAI API. ChatGLM3 was used to validate the effectiveness of our proposed CGPE on LLMs with fewer parameters, while ChatGPT and GPT-4 were utilized to explore the performance ceiling of CGPE without any model training. To ensure the reliability and reproducibility of the experiments, we set the temperature parameter to 0 for all LLMs. In TOG, there are two hyperparameters - path count and exploration steps, both of which were kept at their default values of 3. We modified all prompts in TOG to be suitable for MOOC Q&A, except for entity extraction prompt. The source code of TOG did not include prompt for entity extraction; instead, it directly used entities from the data labels. For a fair comparison, we instructed TOG to use the same entity extraction prompt as CGPE. Additionally, during the experimental process, we observed that TOG sometimes called the LLM several dozen times in a single path exploration, exceeding ten minutes. To mitigate the impact of these extreme cases, we terminated TOG’s exploration when it exceeded 30 LLM calls, providing the explored paths directly to the LLMs for question answering.
Method 1-Hop Judge Multi-Hop Judge 1-Hop Query
partial match complete match
without external knowledge
IO prompt w/CG 8.0 7.3 10.0 7.5
CoT w/CG 3.4 2.2 10.0 6.7
with external knowledge
TOG* w/GLM 34.6 16.2 - -
TOG* w/CG 53.8 39.2 - -
TOG* w/GPT-4 54.3 41.7 - -
TOG w/GLM 5.8 8.1 25.0 12.5
TOG w/CG 26.9 8.3 65.6 46.9
TOG w/GPT-4 27.1 8.3 82.6 56.5
CGPE w/GLM 56.7 50.3 39.6 34.2
CGPE w/CG 85.3 67.4 58.7 50.7
CGPE w/GPT-4 87.2 69.3 81.4 72.9
ChatGPT for answering, ChatGLM3-6B for other steps
TOG 7.7 16.2 28.9 13.8
TOG* 26.9 21.6 - -
CGPE 82.0 63.9 42.3 38.9
Table 1: Results of CGPE for different datasets. CG represents ChatGPT, and GLM represents ChatGLM3-6B.
4.2 Question-Answering
We conducted experiments on three question-answering sets: 1-hop Judge set, multi-hop Judge set, and 1-hop Query. In the Judge set, the knowledge chains corresponding to questions with negative answers are incomplete in the knowledge base. We believe that accurate path exploration should be able to discover incomplete knowledge chains. However, for incomplete knowledge chains, TOG tends to discard them and requires LLMs to directly answer the questions. To facilitate a fair comparison, we modified this logic and designated it as TOG*.
The results are shown in Table 1. It is evident that the baseline methods, IO and CoT, which do not utilize external knowledge, exhibit very low accuracy rates in answering questions. This indicates that MOOC Q&A demands a high reliance on the knowledge base, making it well-suited for evaluating the effectiveness of knowledge base question-answering methods.
On the Judge sets, CGPE demonstrates outstanding performance, surpassing the previous state-of-the-art method TOG. Even the performance of CGPE based on ChatGLM3-6B surpasses that of TOG based on GPT-4. If the answering stage model is switched to ChatGPT, while the other stages use ChatGLM3-6B, CGPE shows a substantial improvement, approaching the effectiveness of methods based on GPT-4. This suggests that, for the Judge set, our proposed CGPE has achieved impressive results in path exploration on ChatGLM3-6B, albeit constrained by the answering capability of ChatGLM3-6B.
On the Query set, considering the partial match metric, TOG based on ChatGPT and GPT-4 performs reasonably well, while its effectiveness diminishes on ChatGLM3-6B. In contrast, our proposed CGPE framework outperforms TOG on ChatGLM3-6B, delivering comparable results to TOG when based on ChatGPT and GPT-4. Examining the exact match metric, CGPE consistently outperforms TOG across all scenarios. This indicates that while TOG can identify partial answers to questions, it struggles to comprehensively retrieve all possible answers to a given question. In contrast, CGPE excels in providing a comprehensive set of answers to the posed questions. Even when substituting the answering-stage model with ChatGPT and utilizing ChatGLM3-6B for other stages, CGPE demonstrates substantial improvement. This underscores the applicability of our proposed method to low-parameter models and emphasizes its practical value.
4.3 Computational Cost
In the knowledge base question-answering process, the majority of computational resources are dedicated to utilizing the Large Language Model, while the remaining computational consumption is relatively negligible. We calculate the average number of LLM calls during the question-answering process to illustrate the computational resource consumption of TOG and CGPE. The results are shown in Table 2. In most cases, TOG has a significantly higher frequency of calls to LLMs, several times more than the calls made by CGPE to LLMs. Additionally, regardless of whether it is dealing with single-hop or multi-hop problems, the frequency of TOG calls remains relatively stable. This is primarily because TOG does not have a clear termination condition; instead, it requires LLMs to determine whether termination is appropriate.
Method 1-Hop Judge Multi-Hop Judge 1-Hop Query
TOG w/ChatGLM3-6B 21.21 24.35 21.6
TOG w/ChatGPT 19.81 22.16 11.2
TOG w/GPT-4 22.6 26.17 8.06
CGPE w/ChatGLM3-6B 6.14 9.4 3.7
CGPE w/ChatGPT 4.76 8.21 4.93
CGPE w/GPT-4 5.24 9.34 4.03
Table 2: Average Number of Calls to LLMs
In contrast, our proposed CGPE significantly reduced the number of calls to large language models compared to TOG, underscoring CGPE’s low computational resource requirements. Furthermore, the invocation frequency of large language models by CGPE varied insignificantly across different backbone models but showed variations across different question-answer sets. This discrepancy arises from the differing knowledge path lengths required to answer questions in distinct question-answer sets. This observation suggests that within the CGPE framework, the termination of exploration based on clues can adaptively adjust the depth of exploration, minimizing the demand for computational resources. Further experiments to delve into the termination of exploration will be conducted in subsequent ablation studies.
Method acc avg LLM Calls
Answering Only 8.0 1
+ LLM Pruning 61.3 8.98
+ Character Matching 82.0 8.02
+ Termination 82.0 6.14
Table 3: Ablation Studies of Matching Step on 1-Hop Judge set. The steps for answering questions are based on ChatGPT, while all other steps are performed using ChatGLM3-6B.
4.4 Ablation Studies
4.4.1 Ablation Study on Path Exploration
CGPE comprises three key stages: clue extraction, path exploration, and answering. Clue extraction serves the path exploration stage, and it cannot be experimented as an independent stage. For answering stage, we adopted a straightforward approach. Thus, we only conducted ablation studies focusing on the path exploration stage. The path exploration stage can be divided into three parts: character matching, large language model pruning, and termination of exploration. The large language model pruning includes fuzzy matching, rough matching and exact matching. The results of ablation studies are shown in Table 3.
We treat direct questioning of ChatGPT as the blank control group, progressively incorporating LLM Pruning, character matching, and termination of exploration to validate the effects of each step. When not using the termination of exploration, we explored each path thrice, obtaining knowledge chains with three nodes. In these experiments, we utilized ChatGPT for answering questions, while all other steps employed ChatGLM. From the table 3, it is evident that solely employing LLM pruning step yields most of the benefits of the complete process, but the invocation of the large language model remains high at 8.98 times.
Introducing character matching not only improves effectiveness but also reduces the frequency of invoking the large language model. Choosing clues, rather than depth of exploration, as the termination criterion has no impact on effectiveness but further reduces the number of invocations of the large language model.
As a summary, the clue-guided LLM pruning step constitutes the majority of the overall stage effectiveness, while the roles of the other two steps primarily involve reducing computational costs.
CE PE A 1-Hop Query
partial match complete match
GLM GLM CG 42.3 38.9
CG GLM CG 47.3 40.7
G4 GLM CG 55.3 49.3
G4 GLM G4 54.7 50.7
GLM GLM GLM 39.6 34.2
CG CG CG 58.7 50.7
G4 G4 G4 80.5 71.4
Table 4: Performances with different backbone models for each stage. PE represents the path exploration stage, CE represents the clue extraction stage, and A represents the answering stage. GLM represents ChatGLM3-6B, CG represents ChatGPT, and G4 represents GPT-4.
4.4.2 Performances With Different Backbone Models
Thanks to the plug-and-play nature of CGPE, we conducted experiments on the three stages of the framework—Clue Extraction, Path Exploration, and Answering—using different large language models with varying capabilities: ChatGLM3-6B, ChatGPT, and GPT-4. The purpose of these experiments is to validate the impact of models with varying capabilities on the effectiveness at each stage. These experiments were conducted on the Query set because, as demonstrated in Table 1, different models have a significant impact on the experimental results. It is conducive to showcasing the diverse influences of various models on each stage.
The results are shown in Table 4. It is evident that substituting a model with higher capabilities at any stage enhances the final result, albeit with varying degrees of improvement across stages. During the answering stage, transitioning from ChatGLM to ChatGPT leads to noticeable improvements. However, the incremental gain is marginal when upgrading from ChatGPT to the more advanced GPT-4, as evidenced by the results on Judge sets in Table 1. This observation suggests that our framework may have a threshold in terms of the model’s capabilities required for the answering stage, where ChatGPT already demonstrates substantial adequacy. In the stages of clue extraction and path exploration, substituting a model with higher capabilities corresponds to superior results, with the most substantial enhancement observed in the path exploration stage. The path exploration stage primarily demands capabilities in information matching. If we fine-tune a less parameterized large language model specifically for “information matching” task Hu et al. (2021), it might achieve satisfactory results. Notably, clue extraction and answering require only one invocation of a large language model each, with the majority of computational costs concentrated in the path exploration stage. A less parameterized large language model for path exploration has the potential to significantly reduce the computational costs of CGPE during actual deployment.
5 Conclusion
In this paper, we proposed a clue-guided path exploration framework named CGPE, designed to assist large language models in answering questions by exploring knowledge paths from knowledge bases. For a given question, CGPE initially extracts clues from the question and then explores knowledge paths guided by these clues. At each node along the knowledge path, we require a match with the extracted clues. Upon matching all clues within a knowledge path, we terminate the exploration of that path and present it to LLMs for answering the question. Experiments conducted on various question-answering sets indicate that CGPE outperforms previous state-of-the-art methods. Moreover, CGPE demonstrates high applicability to low-parameter LLMs compared to existing approaches, making it particularly valuable in scenarios with limited computational resources. Ablation studies validate the effectiveness of each step in the CGPE framework, providing insights for subsequent optimizations. In our future work, we aim to enhance the information matching capability of low-parameter LLMs to achieve better performance with lower computational resource consumption in CGPE.

### v2

@@
-Clue-Guided Path Exploration: An Efficient Knowledge Base Question-Answering Framework with Low Computational Resource Consumption
+Clue-Guided Path Exploration: Optimizing Knowledge Graph Retrieval with Large Language Models to Address the Information Black Box Challenge
 Abstract
@@
 However, updating their knowledge poses challenges, potentially leading to inaccuracies when confronted with unfamiliar queries.
-While integrating knowledge graphs with LLMs has been explored, existing approaches treat LLMs as primary decision-makers, imposing high demands on their capabilities.
-This is particularly unsuitable for LLMs with lower computational costs and relatively poorer performance.
-In this paper, we introduce a Clue-Guided Path Exploration framework (CGPE) that efficiently merges a knowledge base with an LLM, placing less stringent requirements on the model’s capabilities.
-Inspired by the method humans use to manually retrieve knowledge, CGPE employs information from the question as clues to systematically explore the required knowledge path within the knowledge base.
+To address this issue, integrating external knowledge bases such as knowledge graphs with large language models is a viable approach.
+The key challenge lies in extracting the required knowledge from knowledge graphs based on natural language, demanding high semantic understanding.
+Therefore, researchers are considering leveraging large language models directly for knowledge retrieval from these graphs.
+Current efforts typically rely on the comprehensive problem-solving capabilities of large language models.
+We argue that a problem we term the ‘information black box’ can significantly impact the practical effectiveness of such methods.
+Moreover, this kind of methods is less effective for scenarios where the questions are unfamiliar to the large language models.
+In this paper, we propose a Clue-Guided Path Exploration (CGPE) framework to optimize knowledge retrieval based on large language models.
+By addressing the ‘information black box’ issue and employing single-task approaches instead of complex tasks, we have enhanced the accuracy and efficiency of using large language models for retrieving knowledge graphs.
 Experiments on open-source datasets reveal that CGPE outperforms previous methods and is highly applicable to LLMs with fewer parameters.
@@
 For organizations and individuals facing constraints in computational resources, our research offers significant practical value.
-1 Introduction
-Large language models (LLMs) Achiam et al.
-(2023); Ouyang et al.
-(2022); Touvron et al.
-(2023); Cheng et al.
-(2023) have shown their power on many natural language process tasks.
+Introduction
+Large language models (LLMs) (Achiam et al.
+2023; Ouyang et al.
+2022; Touvron et al.
+2023; Cheng, Li, and Bing 2023) have shown their power on many natural language process tasks.
 Nevertheless, there are certain issues associated with LLMs when undertaking question-answering tasks.
 The knowledge of LLMs remains static, frozen at the point of their training, making it challenging to incorporate timely updates.
-Furthermore, when faced with unfamiliar queries, LLMs may sometimes provide incorrect answers with an overconfident tone Luo et al.
-(2023), a phenomenon often referred to as hallucination Bang et al.
-(2023); Ji et al.
-(2023).
-Acknowledging these challenges, an inherent and auspicious resolution involves the integration of external knowledge Mialon et al.
-(2023), specifically employing knowledge graphs (KGs), to enhance the question-answering proficiency of LLMs.
-Figure 1: Comparison among three different types of methods.
-The dashed box indicates the capability requirements of the methods for LLMs.
-The first two types of methods have higher demands on the overall capabilities of LLMs, requiring them to complete comprehensive tasks.
-In contrast, our proposed clue-guided path exploration method has relatively lower requirements on the overall capabilities of LLMs, only demanding them to complete a one task at a time.
+Furthermore, when faced with unfamiliar queries, LLMs may sometimes provide incorrect answers with an overconfident tone (Luo et al.
+2023), a phenomenon often referred to as hallucination (Bang et al.
+2023; Ji et al.
+2023).
+Acknowledging these challenges, an inherent and auspicious resolution involves the integration of external knowledge (Mialon et al.
+2023), specifically employing knowledge graphs (KGs), to enhance the question-answering proficiency of LLMs.
 Researchers have conducted some studies to combine external knowledge and LLMs.
-Some researchers prompt LLMs to generate database operation instructions to retrieve information Hu et al.
-(2023); Wang et al.
-(2023).
-Chain-of-thought Wei et al.
-(2022b) is also used to help LLMs decompose the input question into multiple sub-questions Li et al.
-(2023a); Zhao et al.
-(2023).
+Prompting LLMs to directly generate corresponding knowledge base operation statements based on questions is an intuitive approach (Hu et al.
+2023; Wang et al.
+2023).
+Chain-of-thought (Wei et al.
+2022b) is also used to help LLMs decompose the input question into multiple sub-questions (Li et al.
+2023a; Zhao et al.
+2023a).
 However, in these methods, LLMs lack awareness of the nature of external knowledge sources.
 As a result, the instructions or sub-questions generated by LLMs may not align well with the external knowledge.
-TOG Sun et al.
-(2023b) addresses the mismatch issue between LLMs and knowledge bases by facilitating multi-turn communications, continually supplying updated knowledge path information to LLMs.
-Nevertheless, in the course of TOG’s path exploration, critical decisions, such as choosing relationships, selecting entities, and determining termination points, heavily depend on the inherent capabilities of the large language model.
-TOG lacks the inclusion of any mechanism or provision of supplementary information to aid LLMs in decision-making.
-This presents a significant challenge to the overall capabilities of LLMs and impedes the accurate exploration of knowledge paths.
-As a result, it makes the approach less suitable for LLMs with comparatively weaker comprehensive abilities.
-This issue persists and remains a common challenge faced by previous methods.
-As the capabilities of LLMs strengthen, their parameter sizes increase, necessitating more computational resources, such as GPUs.
-For a substantial proportion of organizations and individuals, the availability of computational resources is a crucial consideration.
-Therefore, developing a mechanism to lower the demands on LLMs’ capabilities during path exploration not only enhances the effectiveness of knowledge base question-answering but also carries significant practical value.
-In everyday scenarios, when confronted with the task of retrieving information from a knowledge base, such as a book, in response to a question, we typically initiate the search process by distilling pertinent details from the inquiry.
-These extracted fragments serve as clues, enabling us to methodically narrow down the search scope and ultimately arrive at the desired answer.
-Each step in this process is straightforward yet remarkably effective.
-Drawing inspiration from this practical approach, we have developed a Clue-Guided Path Exploration (CGPE) framework.
-This framework harnesses the information embedded in questions as clues, guiding a systematic exploration of the knowledge graph and progressively revealing the path to the requisite knowledge.
-For a given question, our methodology initially extracts ‘clues’ from the question.
-Based on these clues, we identify initial nodes of knowledge paths within the knowledge base.
-With the initial nodes, we can retrieve the related nodes.
-Then we match the information of these related nodes with clues to determine the next node for each path.
-We iterate through this process until all clues have undergone a matching cycle.
-We iterate through this process until we find a path that matches all the clues.
-This refined knowledge path is then presented to LLMs for answering the question.
-Notably, our approach positions the clues as guiding indicators throughout the entire path exploration process.
-In doing so, LLMs only need to focus on information matching using the given clues and information, without the added complexity of analyzing questions at the same time.
-This minimizes the requirement for LLMs to possess extensive capabilities, thereby improving the efficiency of path exploration and making our approach suitable for LLMs with fewer parameters.
+To address this mismatch, some researchers have attempted an verify-and-edit approach (Zhao et al.
+2023b; Cheng et al.
+2024), no longer requiring LLMs to directly generate database operation statements.
+Instead, they first generate potential knowledge paths required for questions based on LLMs’ own knowledge, and then edit these knowledge paths based on the graph.
+In the scenario envisioned by these methods, LLMs are assumed to have a complete understanding of the questions and possess relevant knowledge to generate appropriate knowledge paths based on those questions.
+However, it is worth noting that the primary motivation for integrating external knowledge to answer questions lies in the fact that LLMs encounter unfamiliar questions, especially in some specialized domains.
+This conflicts with the ideal scenario for such methods.
+To make the method more applicable to unfamiliar questions, researchers no longer require LLMs to ”imagine” answers out of thin air.
+Instead, they provide information from the graph to LLMs, allowing them to select appropriate information until they have gathered enough to answer the question (Sun et al.
+2023b).
+This approach reduces the demand for the LLMs’ own knowledge and emphasizes the utilization of their comprehensive capabilities.
+Ideally, an LLM with sufficiently powerful comprehensive capabilities can analyze the question and progressively select corresponding knowledge from the graph without repetition.
+However, on one hand, current LLMs are still insufficiently capable of perfectly performing these tasks, which limits the practical effectiveness of this approach.
+On the other hand, greater capabilities often necessitate larger parameter sizes, which in turn increases the demand for computational resources.
+This poses a challenge to the widespread application of this method.
+To achieve better performance with lower computational resource requirements, it is essential to research how to reduce the demands on LLMs.
+We believe that existing methods encounter what we refer to as the ”information black box” problem.
+In current approaches, LLMs handle all tasks related to information analysis and processing.
+During this process, it is unclear what essential information the LLMs extract from the questions or which pieces of information are used when exploring the knowledge graph.
+Given that the capabilities of LLMs are not yet sufficient to flawlessly perform all tasks, some information may be overlooked or redundantly used, leading to potential gaps or repetitions in the exploration of the knowledge graph.
+In this paper, we propose a clue-guided path exploration framework called CGPE, which addresses the information black box problem while reducing the demands on the capabilities of LLMs, enabling more efficient exploration of knowledge graph.
+Unlike existing methods that analyze and use information in a black-box manner, CGPE explicitly extracts all necessary information from the questions.
+This information is treated as clues and guides the entire path exploration process.
+We utilize “information matching” tasks to implement each step of CGPE’s path exploration process, transforming the requirement from comprehensive capabilities of LLMs to a singular demand for “information matching”.
+It is noteworthy that both extracting and matching information do not necessitate a deep understanding by the LLMs, since other information can also assist in these tasks.
+This aids LLMs in tackling “unfamiliar problems.” Detailed discussion on this will be provided in Chapter Three.
+Our approach offers a promising direction for LLMs to explore knowledge graphs, showcasing significant practical implications.
+Figure 1: Comparison between CGPE and Existing Methods.
 The main contributions of this paper are summarized as follows:
-  • We present CGPE, a novel clue-guided path exploration framework aimed at assisting large language models (LLMs) in answering questions.
-By incorporating a path exploration mechanism centered around clues, we have effectively lower the demands on LLMs’ capabilities.
-  • Experiments conducted on various datasets demonstrate that our proposed method outperforms the previous state-of-the-art method.
-Notably, our approach performs well on LLMs with fewer parameters, marking a significant breakthrough compared to previous methods.
-  • Experiments show that our method involves very few invocations of LLMs, significantly reducing computational resource consumption.
-Ablation studies offer insights for further reducing the consumption of computational resources.
-2 Related Work
-2.1 Knowledge Enhanced Large Language Models
-One direct approach to addressing the generation of illusions by LLMs is the use of external knowledgeMialon et al.
-(2023) to enhance the model’s ability to perceive the real world.
+  • To the best of our knowledge, we are the first to identify the “information black box” problem in current methods.
+By addressing this problem, we aim to utilize the information from questions more clearly and efficiently, thereby enhancing the accuracy and efficiency of Large Language Models (LLMs) in retrieving knowledge from knowledge graphs.
+  • We present CGPE, a novel clue-guided path exploration framework aimed at assisting LLMs in retrieving knowledge from knowledge graphs.
+Unlike existing methods that center around LLMs, our framework is entirely designed around the information contained within the questions.
+This provides a new paradigm for future research on LLMs in retrieving knowledge from knowledge graphs.
+  • We employed a single “information matching” task to explore knowledge paths in knowledge graphs.
+Unlike complex tasks in existing methods, this singular task shows promising results even with fewer parameters in LLMs, enhancing practical utility.
+Moreover, fine-tuning LLMs based on this single task may further enhance the effectiveness of our approach.
+  • Multiple sets of experiments have demonstrated that our proposed CGPE framework offers various advantages over existing methods in knowledge graph retrieval.
+These include higher accuracy, reduced frequency of calls to LLMs, and effective performance even with fewer parameters in LLMs.
+These advantages underscore the practical value of our approach.
+Related Work
+Knowledge Enhanced Large Language Models
+One direct approach to addressing the generation of illusions by LLMs is the use of external knowledge (Mialon et al.
+2023) to enhance the model’s ability to perceive the real world.
 This has led to the emergence of numerous knowledge-based models.
-Information retrieval is commonly employed in developing knowledge-based chatbotsShuster et al.
-(2021).
-BlenderBot2Xu et al.
-(2021) integrates Internet search capabilities.
-SeeKeRShuster et al.
-(2022a) surpasses BlenderBot2Xu et al.
-(2021) by executing three modular tasks using a single language model: generating search queries, extracting relevant knowledge from retrieved documents, and generating the final response.
-BlenderBot3Shuster et al.
-(2022b) fine-tunes a 175-billion-parameter OPTZhang et al.
-(2022) on a combination of 20 question-answering and dialogue datasets.
-AtlasIzacard et al.
-(2023) stands out as a top model on the KILT benchmarkPetroni et al.
-(2021), which includes 11 knowledge-oriented tasks, including Wizard of WikipediaDinan et al.
-(2019).
-2.2 Knowledge base Enhanced Models
+Information retrieval is commonly employed in developing knowledge-based chatbots (Shuster et al.
+2021).
+BlenderBot2 (Xu, Szlam, and Weston 2021) integrates Internet search capabilities.
+SeeKeR (Shuster et al.
+2022a) surpasses BlenderBot2 (Xu, Szlam, and Weston 2021) by executing three modular tasks using a single language model: generating search queries, extracting relevant knowledge from retrieved documents, and generating the final response.
+BlenderBot3 (Shuster et al.
+2022b) fine-tunes a 175-billion-parameter OPT (Zhang et al.
+2022) on a combination of 20 question-answering and dialogue datasets.
+Atlas (Izacard et al.
+2023) stands out as a top model on the KILT benchmark (Petroni et al.
+2021), which includes 11 knowledge-oriented tasks, including Wizard of Wikipedia (Dinan et al.
+2019).
+Knowledge base Enhanced Models
 There are diverse ways to enhance models based on knowledge, aiming to address issues arising from complex reasoning, hallucination, and poor interpretability.
-Cao et alCao et al.
-(2021) improved implicit event causality extraction by leveraging structured knowledge bases.
-Zhao et alZhao et al.
-(2023) recently proposed a Verification and Editing (VE) framework.
-This framework improves the Chain of Thought (CoT) reasoning in large language models (LLMs) by integrating retrieval systemsWei et al.
-(2022a).
-Researchers have also introduced the Chain-of-Knowledge (CoK)Li et al.
-(2023b), a novel framework designed to enhance the factual correctness of LLMs using heterogeneous knowledge sources.
+Cao et al (Cao et al.
+2021) improved implicit event causality extraction by leveraging structured knowledge bases.
+Zhao et al (Zhao et al.
+2023a) recently proposed a Verification and Editing (VE) framework.
+This framework improves the Chain of Thought (CoT) reasoning in large language models (LLMs) by integrating retrieval systems (Wei et al.
+2022a).
+Researchers have also introduced the Chain-of-Knowledge (CoK) (Li et al.
+2023b), a novel framework designed to enhance the factual correctness of LLMs using heterogeneous knowledge sources.
 This ensures that inaccuracies in the foundational principles do not propagate to subsequent steps.
-Think-on-Graph(ToG) Sun et al.
-(2023a) is a plug-and-play framework that seamlessly applies to various LLMs and KGs.
-ToG also enhances the reasoning capabilities, enabling smaller LLMs (such as LLAMA2-70B) to compete with larger LLMs (such as GPT-4).
-3 Method
-As shown in Figure 2, our framework can be divided into three stages: clue extraction stage, path exploration stage, and answering stage.
-The clue extraction stage is elaborated in Section 3.2, while the path exploration stage is discussed in Sections 3.3, 3.4, and 3.5.
-As for the answering stage, we did not implement any specific design but rather provided the extracted knowledge to Large Language Models in the form of triplets to assist them in answering questions.
-Figure 2: The workflow of our proposed CGPE framework.
-The green box indicates that these steps require the use of LLMs.
-The process of exploring the next node is within the dashed box, and it continues until the termination condition is met.
-3.1 Problem Formulation
-For a given question, we initially extract a set of clues, denoted as $C=\{c_{1},c_{2},\ldots,c_{n}\}$.
-We employed $P=\{p_{1},p_{2},\ldots,p_{m}\}$ to denote all paths explored within the knowledge base, where each $p_{i}$ represents an individual path encompassing all entities along that path.
-Each node potentially corresponds to a clue $c_{i}$ from set $C$.
-The clues corresponding to the nodes are marked as ‘used’ for the path, and they will not be used again to explore this path.
-Our objective is to identify a path that corresponds to all clues $c_{i}$ within set $C$.
-Ultimately, we present this curated path to the LLM to facilitate the answering of the given question.
-3.2 Extracting Clues from the Given Question
+Think-on-Graph(ToG) (Sun et al.
+2023a) is a plug-and-play framework that seamlessly applies to various LLMs and KGs.
+ToG also enhances the reasoning capabilities, enabling smaller LLMs to compete with larger LLMs.
+Method
+To enable more efficient exploration of knowledge graphs by large language models (LLMs) under lower computational resource constraints, CGPE addresses the information black box problem inherent in existing methods through a clear information utilization mechanism.
+This mechanism is executed using a single task to perform each step, thereby reducing the comprehensive capabilities required from the LLMs.
+Necessary Information
+Firstly, we need to clarify which pieces of information are required for our purposes.
 Intuitively, we posit that the majority of information (not all information) within a question can be characterized by entities and relationships originating from the question.
-Particularly, the knowledge required to answer a question is presumed to be related with these entities and relationships.
 Employing these entities and relationships as clues for knowledge retrieval aids in addressing the given question.
-For instance, in the case of the question “What is the population of the capital of China?” utilizing (China, capital, population) as clues facilitates the effortless retrieval of relevant knowledge to answer the question.
-The delineation between entities and relationships is not always clear-cut.
-In the aforementioned example, ‘capital’ can function both as a relationship and an entity.
-Additionally, relationships can, to some extent, be represented and inferred by entities Bordes et al.
-(2013); Sun et al.
-(2019).
-For instance, with an incomplete triple (China, ?, Beijing), one can infer the missing relationship as ‘capital’.
-Consequently, in our practical extraction approach, we do not rigidly distinguish between entities and relationships.
-Instead, we broadly task the LLM with extracting all entities from the question, utilizing them as clues to guide the exploration of paths within the knowledge graph.
-In this scenario, words that can function both as entities and relationships are also extracted.
-3.3 Selection of Initial Nodes
-Upon extracting clues from a given question, we initiate a path exploration within the knowledge graph.
-To choose the initial nodes for the paths, we conduct node queries in the knowledge graph using clues within the set $C$ as the value of nodes.
-In the scenario where the knowledge graph contains the necessary information to answer questions, we typically achieve successful queries for a specific clue (i.e.
-node).
-there invariably exists certain clue $C_{i}$ that can be retrieved, corresponding to specific nodes.
-These nodes serve as the initial nodes for paths, with the number of paths aligning with the quantity of nodes.
-And the corresponding clues are marked as ‘used’ for each path.
-It is important to note that although these initial nodes mark the inception of a path, they may not necessarily be the commencement of the target knowledge chain; they could also occupy an intermediary position within the desired knowledge chain.
-This may pose a challenge for path exploration.
-We will address this issue in subsequent steps.
-3.4 Exploration of the Next Node
-For each path, we iteratively explore the next node until the nodes along the path collectively cover all clues in set $C$.
-Employing a predefined query for retrieving related nodes of the current node, we obtain a candidate set of these related nodes, capturing relationships, entities, and entity labels.
-The labels serve as abstract summaries of entity information, aligning with the abstract and non-specific nature in which clues are often presented in questions.
-Most knowledge graphs label entities, and if a label is absent, it is treated as ‘none’.
-Subsequently, we proceed through a four-step process to match the candidate set with clues.
-Character Matching.
-Just as discussed when selecting the initial node, the initial node for a path is matched with the characters of a certain clue.
-If a related node matches an unused clue, then we consider this node as the next one in the path.
-And the clue $c_{i}$ is marked as ‘used’ for the path.
-Fuzzy Matching.
-If no node is selected after character matching, we then prompt the LLM to conduct a fuzzing matching.
-Nodes in the candidate set are characterized with (relationship, label), which is a highly abstract representation of node information.
-We prompt LLM to match (relationship, label) with unused clues.
-If the model deems (relationship, label) as unrelated to any unused clue, the corresponding entities are pruned from the candidate set.
-Note that different nodes may be characterized with the same (relationship, label).
-Rough Matching.
-The remaining nodes in the candidate set are then represented as (label, entity), in which the label is used to provide abstract information and the entity is used to provide specific information about the node.
-We prompt LLM to match and score the (label, entity) with unused clues.
-Considering that non-fine-tuned LLMs may not provide highly accurate scores, we simply remove nodes with a score of 0.
-But with a fine-tuned LLM, selecting the top-k nodes can reduce the frequency of invoking the LLM during the exact matching step.
-Exact Matching.
-We perform exact matching on the remaining nodes to ultimately determine the next node in the path.
-The remaining nodes are still represented as (label, entity), and we prompt LLM to match each (label, entity) with unused clues.
-For each (label, entity), the most compatible clue $c_{i}$ is marked as ‘used’, and the corresponding node is chosen as the next node.
-If the LLM does not find a matching clue for the (label, entity), then the corresponding node is discarded.
-If multiple nodes are chosen as the next node, this step generates multiple new paths.
-Considering the scenario where initial nodes may lie within the middle of the target knowledge chain, theoretically, two paths exploring both ends of the target knowledge chain may be generated.
-The rough matching and exact matching steps share similarities in terms of the information provided to the LLM.
-The distinction lies in their respective emphases, with the former prioritizing screening and the latter emphasizing precision.
-This division is deemed advantageous for achieving favorable outcomes on relatively less capable LLMs.
-Given the relatively low computational cost of these LLMs, the trade-off between multiple computations and enhanced effectiveness is deemed worthwhile.
-For LLMs with high capabilities and significant computational demands, keeping only the rough matching step is also acceptable.
-3.5 Termination of Exploration
-If a path corresponds to all clues $c_{i}$ within set $C$, it is considered to encompass all the necessary knowledge to answer the question.
-We then terminate the exploration of this path and use it as a knowledge chain to help the LLM answer the given question.
-In instances where paths still possess unused clues, yet fail to identify the next nodes, exploration for those paths is also terminated.
-Such paths can be categorized into two distinct types.
-The first is an incorrect path that ought to be disregarded.
-The second entails the correct path, wherein the initial node of the path is situated at the midpoint of the target knowledge chain, thereby leading to the exploration of nodes in a unidirectional manner.
-As previously discussed, another path theoretically exists, exploring the opposite direction.
-Merging these two paths usually ensures coverage of all clues in set $C$.
-Upon the discovery of a path containing the requisite knowledge to answer the question, exploration of other paths is halted.
-The identified path is then presented to the LLM in the form of triplets, helping the LLM to answer the given question.
-4 Experiments
-4.1 Experimental Settings
+For instance, in the question ”What is the population of the capital of China?” the key information required is ”population,” while ”China” and ”capital” serve to narrow down the context.
+Utilizing the clues (China, capital, population) can help us retrieve the correct answer from the knowledge graph, but omitting any one of these elements may lead to an incorrect answer.
+However, when faced with unfamiliar questions, such as those in specialized domains, LLMs—especially those with limited computational resources—may struggle to accurately extract entities and relationships from the questions posed.
+It’s worth noting that extracting entities from sentences doesn’t necessarily require understanding those entities; other information in the sentence, including syntactic structures, can aid in extracting unfamiliar entities.
+However, extracting relationships between two entities, especially implicit relationships, often requires a thorough understanding of the entities, which can be challenging when facing unfamiliar questions.
+On the other hand, in knowledge graphs, there typically aren’t multiple different relationships between the same pair of entities.
+Once the entities in a knowledge path are determined, the relationships often follow suit.
+Taking these considerations into account, we prompt LLMs to extract the entities in the question as the clues to guide the subsequent path exploration.
+Clues to Knowledge Paths
+Now that we have the necessary clues $C=\{c_{1},c_{2},\ldots,c_{n}\}$, and we assume that clues derived from a logically sound question should be interconnected, mirroring their corresponding knowledge nodes in a knowledge graph, thus forming a continuous knowledge path.
+Our goal is to identify a knowledge path $P=\{e_{1},r_{1,2},e_{2},r_{2,3},e_{3},\ldots,e_{n}\}$ corresponding to these clues, such that $P$ contains the knowledge required to answer the question.
+The $e_{i}$ in $P$ represents an entity in the knowledge graph, and the $r_{i,j}$ represents a relation between $e_{i}$ and $e_{j}$.
+It is important to note that although we only extract entities from the question, our clues may correspond to a relationship in the graph rather than an entity.
+For such a clue $c_{i}$ and the corresponding relation $r_{k,j}$, the nodes $e_{j}$ related to $r_{k,j}$ should also relate to $c_{i}$.
+To eliminate the uncertainties, we assume that each clue corresponds to the nodes during the subsequent path exploration process.
+Path Exploration
+To start the path exploration, we need to find an initial point.
+For any $c_{i}$, if we can retrieve the corresponding node $e_{i}$ from the graph through character matching, then $e_{i}$ can be considered as the initial point.
+For CGPE, identifying just one such initial point is sufficient for subsequent path exploration.
+Even the initial point may be a node in the middle of the path $P$, rather than one of its endpoints.
+Existing methods either assume paths generated by LLMs align closely with the graph (Zhao et al.
+2023b; Cheng et al.
+2024) or treat the endpoints of paths as given conditions (Sun et al.
+2023a).
+In contrast, our method imposes fewer usage constraints.
+Typically, such initial points are readily discoverable if the relevant knowledge exists in the graph, as confirmed by our experimental results.
+Next, we iteratively explore each subsequent node until the entire path is obtained.
+Assuming the current node is $e_{i}$, we retrieve all associated edges and neighboring nodes of $e_{i}$ from the graph, and then employ a three-step matching process to identify a node that can be matched with any unused clue.
+That node then be added into $P$ and become the “current node” for the next round of exploration.
+To systematically utilize clues without redundancy, we adhere to several rules:
+  • Rule 1, previously used clues are excluded from subsequent matches;
+  • Rule 2, if multiple nodes successfully match during the same round of exploration, each node branches into a candidate path for independent exploration;
+  • Rule 3, if there exists a path $P$ that matches all clues, or multiple paths can be combined to form such a $P$, then $P$ serves as the knowledge provided to LLMs for answering questions.
+The rule 3 ensures that even if the initial point is an intermediate node of $P$, the path exploration can proceed correctly.
+Information Matching
+To reduce the dependency on the comprehensive capabilities of the LLMs, we employ a single “information matching” task to complete the aforementioned three-step matching process.
+To complete an information matching task, it is not necessary to be fully familiar with the entities involved; a general understanding of their labels, categories, and similar attributes suffices.
+This is highly beneficial for addressing unfamiliar questions.
+Given the often excessive number of entities, a preliminary filtering based on summary information about the entities can enhance matching efficiency.
+In the first step, We retrieve from the graph all entities related to the current entity $e_{i}$ as candidates, where $e_{i}$ is the last node in path $P$.
+The summary information of all candidate entities, such as (relation, label), is used to represents the candidate entities.
+We then prompt LLMs to match the summary information with the unused clues.
+When the summary information successfully matches with any cue, the entity associated with that summary information progresses to the subsequent stage of matching.
+Since the summary information may be identical for multiple distinct entities, utilizing summary information for matching can reduce the number of matching attempts and enhance overall matching efficiency.
+For the second step, candidate entities are represented using a combination of summary information and specific details, such as (label, entity name).
+We then prompt the LLMs to match these representations with the unused clues.
+When the information successfully matches with any cue, the entity associated with that information progresses to the subsequent stage of matching.
+Following the initial coarse filtering step, this stage involves a more comprehensive evaluation to further refine the candidate entities based on additional detailed information.
+In the third step, the candidate entities are still represented using a combination of summary information and specific details.
+Here, the LLMs independently matches them with clues and specifies which clue successfully matched each entity.
+This step establishes a direct association between entities and specific clues.
+After completing the three-step matching process, the resulting node $e_{i+1}$ and the relation $r_{i,i+1}$ between $e_{i+1}$ and $e_{i}$ are added to path $P$.
+Note that if any unused clue $c_{i}$ can match any candidate node from the first step through character matching, then it should be added to path $P$.
+We then check whether there exists a path $P$ that satisfies Rule 3.
+If such a path is found, the path exploration process is concluded.
+This termination method helps to reduce the number of explorations and enhance exploration efficiency.
+Experiments
+Experimental Settings
+Datasets and Evaluation Metrics.
+To evaluate the performance of our proposed CGPE in answering questions including familiar and unfamiliar questions, we selected two distinct knowledge graphs as external knowledge bases.
+The first is the knowledge graph from MOOC Q&A (Yu et al.
+2020), and the second is Freebase (Bollacker et al.
+2008).
+MOOC Q&A is a dataset collected from MOOC websites, comprising a knowledge graph and associated question-answer sets.
+The knowledge graph includes 7 categories of entities, totaling 15,388, with 11 types of relationships forming 277,487 triplets.
+The question-answer sets consist of non-trivial questions in the education domain, such as ”Which courses does Professor A teach?” These types of questions are unfamiliar to LLMs.
+The question-answer sets contain Judge set, Query set and Count set.
+Among them, the questions in the Judge Set are all judgment-based, involving determining correctness or incorrectness.
+The questions in both the Query Set and Count Set are similar, involving one-to-many relationships where the answers include multiple entities.
+The difference lies in the fact that the Count Set requires a significantly larger number of entities in the answer compared to the former.
+Freebase is a large-scale, semi-structured database supported by Google, designed to collect and connect information about millions of entities and their relationships worldwide.
+It covers a wide range of topics such as people, places, books, movies, music, etc., with each entity having a unique identifier.
+We evaluate CGPE on two QA sets with Freebase as external knowledge base: WebQSP (Yih et al.
+2016) and WebQuestions (Berant et al.
+2013).
 […]
 Baselines.
-We compared our approach with three baseline methods: standard prompting (IO prompt) Brown et al.
-(2020), chain of thought prompting (CoT prompt) Wei et al.
-(2022b) and TOG Sun et al.
-(2023b).
+We compared our approach with three baseline methods: standard prompting (IO prompt) (Brown et al.
+2020), chain of thought prompting (CoT prompt) (Wei et al.
+2022b) and ToG (Sun et al.
+2023b).
 The knowledge base and Q&A sets in MOOC Q&A are highly interconnected, with most questions requiring knowledge retrieval from the knowledge base for answers.
@@
 Experiment Details.
-We conducted experiments on three different LLMs, namely ChatGLM3 Du et al.
-(2022) with 6 billion parameters, ChatGPT, and GPT-4.
+For MOOC Q&A, we conducted experiments on three different LLMs, namely ChatGLM3-6B (Du et al.
+2022), ChatGPT, and GPT-4.
 ChatGPT and GPT-4 are called by OpenAI API.
-ChatGLM3 was used to validate the effectiveness of our proposed CGPE on LLMs with fewer parameters, while ChatGPT and GPT-4 were utilized to explore the performance ceiling of CGPE without any model training.
+ChatGLM3-6B was used to validate the effectiveness of our proposed CGPE on LLMs with fewer parameters, while ChatGPT and GPT-4 were utilized to explore the performance ceiling of CGPE without any model training.
+For WebQSP and WebQuestions, we conducted experiments on three different LLMs: GLM4-9B-Chat (GLM et al.
+2024) with 9 billion parameters, Llama3-8B-Instruct (Meta 2024).
+Llama is not good at Chinese, so we did not test it on MOOC Q&A.
+Due to economic considerations and OpenAI’s policies, we did not test ChatGPT and GPT-4 for WebQSP and WebQuestions.
 To ensure the reliability and reproducibility of the experiments, we set the temperature parameter to 0 for all LLMs.
@@
 We modified all prompts in ToG to be suitable for MOOC Q&A, except for entity extraction prompt.
-The source code of TOG did not include prompt for entity extraction; instead, it directly used entities from the data labels.
+The source code of ToG does not include prompt for entity extraction; instead, it regards the topic entities as given conditions.
 For a fair comparison, we instructed ToG to use the same entity extraction prompt as CGPE.
@@
 To mitigate the impact of these extreme cases, we terminated ToG’s exploration when it exceeded 30 LLM calls, providing the explored paths directly to the LLMs for question answering.
+Table 1: Results of CGPE for MOOC Q&A. Questions here are unfamiliar to LLMs.
 Method 1-Hop Judge Multi-Hop Judge 1-Hop Query
@@
 without external knowledge
-IO prompt w/CG 8.0 7.3 10.0 7.5
-CoT w/CG 3.4 2.2 10.0 6.7
+IO prompt w/ChatGPT 8.0 7.3 10.0 7.5
+CoT w/ChatGPT 3.4 2.2 10.0 6.7
 with external knowledge
-TOG* w/GLM 34.6 16.2 - -
-TOG* w/CG 53.8 39.2 - -
+ToG* w/GLM3-6B 34.6 16.2 - -
+ToG* w/ChatGPT 53.8 39.2 - -
 ToG* w/GPT-4 54.3 41.7 - -
-TOG w/GLM 5.8 8.1 25.0 12.5
-TOG w/CG 26.9 8.3 65.6 46.9
+ToG w/GLM3-6B 5.8 8.1 25.0 12.5
+ToG w/ChatGPT 26.9 8.3 65.6 46.9
 ToG w/GPT-4 27.1 8.3 82.6 56.5
-CGPE w/GLM 56.7 50.3 39.6 34.2
-CGPE w/CG 85.3 67.4 58.7 50.7
+CGPE w/GLM3-6B 56.7 50.3 39.6 34.2
+CGPE w/ChatGPT 85.3 67.4 58.7 50.7
 CGPE w/GPT-4 87.2 69.3 81.4 72.9
@@
 CGPE 82.0 63.9 42.3 38.9
-Table 1: Results of CGPE for different datasets. CG represents ChatGPT, and GLM represents ChatGLM3-6B.
-4.2 Question-Answering
-We conducted experiments on three question-answering sets: 1-hop Judge set, multi-hop Judge set, and 1-hop Query.
-In the Judge set, the knowledge chains corresponding to questions with negative answers are incomplete in the knowledge base.
-We believe that accurate path exploration should be able to discover incomplete knowledge chains.
-However, for incomplete knowledge chains, TOG tends to discard them and requires LLMs to directly answer the questions.
+Performance Comparison
+Table 2: Results of CGPE for WebQuestions and WebQSP.
+Questions are familiar to LLMs.
+To more clearly demonstrate the ability of ToG and CGPE to retrieve knowledge from the knowledge graph, we require LLMs to answer questions based only on the retrieved knowledge, resulting in the outcomes of ToG - IO and CGPE - IO respectively.
+Method WebQuestions WebQSP
+partial match complete match partial match complete match
+without external knowledge
+IO prompt w/Llama3-8B-Instruct 44.4 37.2 52.4 27.0
+IO prompt w/GLM4-9B-Chat 52.4 33.4 60.1 31.8
+COT w/Llama3-8B-Instruct 54.3 35.8 62.1 33.7
+COT w/GLM4-9B-Chat 54.8 34.3 63.0 32.5
+with external knowledge
+ToG - IO w/Llama3-8B-Instruct 55.1 36.7 55.6 32.3
+ToG w/Llama3-8B-Instruct 67.3 43.9 69.7 39.4
+ToG - IO w/GLM4-9B-Chat 63.3 37.8 63.3 30.6
+ToG w/GLM4-9B-Chat 71.4 46.9 74.0 38.8
+CGPE - IO w/Llama3-8B-Instruct 67.3 42.9 60.8 34.2
+CGPE w/Llama3-8B-Instruct 72.4 45.9 75.0 43.3
+CGPE - IO w/GLM4-9B-Chat 67.4 38.9 62.5 35.7
+CGPE w/GLM4-9B-Chat 78.9 50.5 78.6 48.2
+We conducted experiments on three question-answering sets of MOOC Q&A: 1-hop Judge set, multi-hop Judge set, and 1-hop Query.
+In the Judge set, the knowledge paths corresponding to questions with negative answers are incomplete in the knowledge base.
+However, for incomplete knowledge paths, ToG tends to discard them and requires LLMs to directly answer the questions.
 To facilitate a fair comparison, we modified this logic and designated it as ToG*.
-The results are shown in Table 1.
+The results are shown in 1.
 It is evident that the baseline methods, IO and CoT, which do not utilize external knowledge, exhibit very low accuracy rates in answering questions.
-This indicates that MOOC Q&A demands a high reliance on the knowledge base, making it well-suited for evaluating the effectiveness of knowledge base question-answering methods.
+This implies that these questions are not familiar to LLMs, and answering them requires a high degree of reliance on knowledge within the knowledge graph.
 On the Judge sets, CGPE demonstrates outstanding performance, surpassing the previous state-of-the-art method ToG.
 Even the performance of CGPE based on ChatGLM3-6B surpasses that of ToG based on GPT-4.
-If the answering stage model is switched to ChatGPT, while the other stages use ChatGLM3-6B, CGPE shows a substantial improvement, approaching the effectiveness of methods based on GPT-4.
+If the answering stage model is switched to ChatGPT, while the path exploration stage uses ChatGLM3-6B, CGPE shows a substantial improvement, approaching the effectiveness of methods based on GPT-4.
 This suggests that, for the Judge set, our proposed CGPE has achieved impressive results in path exploration on ChatGLM3-6B, albeit constrained by the answering capability of ChatGLM3-6B.
+This fully aligns with our expectations: CGPE is specifically designed for path exploration in low-computing resource scenarios to enhance performance.
 On the Query set, considering the partial match metric, ToG based on ChatGPT and GPT-4 performs reasonably well, while its effectiveness diminishes on ChatGLM3-6B.
@@
 In contrast, CGPE excels in providing a comprehensive set of answers to the posed questions.
-Even when substituting the answering-stage model with ChatGPT and utilizing ChatGLM3-6B for other stages, CGPE demonstrates substantial improvement.
-This underscores the applicability of our proposed method to low-parameter models and emphasizes its practical value.
-4.3 Computational Cost
+The results of WebQuestions and WebQSP are shown in 2.
+Due to economic considerations and OpenAI’s policies, we did not test ChatGPT and GPT-4 for WebQSP and WebQuestions.
+From the baseline, we can observe that LLMs can answer a considerable portion of the questions solely based on their own knowledge, indicating that these questions are relatively familiar to LLMs.
+To more clearly demonstrate the ability of ToG and CGPE to retrieve knowledge from the knowledge graph, we require LLMs to answer questions based only on the retrieved knowledge, resulting in the outcomes of ToG - IO and CGPE - IO respectively.
+However, even with this requirement, the final results may still be relatively inflated due to the influence of the LLMs’ inherent knowledge.
+It can be seen that for relatively familiar questions, the results of CGPE are still significantly higher than those of ToG.
+This again demonstrates the effectiveness of our proposed method.
+The CGPE - IO results are better than those of ToG - IO, but not as much as the advantage of CGPE over ToG.
+This suggests that the knowledge retrieved by CGPE - IO is more unfamiliar to LLMs, while the knowledge retrieved by ToG - IO is more familiar to LLMs, which also reflects the effectiveness of our method for unfamiliar questions.
+Computational Cost
 In the knowledge base question-answering process, the majority of computational resources are dedicated to utilizing the Large Language Model, while the remaining computational consumption is relatively negligible.
@@
 This is primarily because ToG does not have a clear termination condition; instead, it requires LLMs to determine whether termination is appropriate.
+Table 3: Average Number of Calls to LLMs
 Method 1-Hop Judge Multi-Hop Judge 1-Hop Query
@@
 CGPE w/GPT-4 5.24 9.34 4.03
-Table 2: Average Number of Calls to LLMs
 In contrast, our proposed CGPE significantly reduced the number of calls to large language models compared to ToG, underscoring CGPE’s low computational resource requirements.
@@
 Further experiments to delve into the termination of exploration will be conducted in subsequent ablation studies.
+Table 4: Ablation Studies of Matching Step on 1-Hop Judge set. The answering step is based on ChatGPT, while all other steps are performed using ChatGLM3-6B.
 Method acc avg LLM Calls
 Answering Only 8.0 1
-+ LLM Pruning 61.3 8.98
++ Three Step Matching 61.3 8.98
 + Character Matching 82.0 8.02
 + Termination 82.0 6.14
-Table 3: Ablation Studies of Matching Step on 1-Hop Judge set. The steps for answering questions are based on ChatGPT, while all other steps are performed using ChatGLM3-6B.
-4.4 Ablation Studies
-4.4.1 Ablation Study on Path Exploration
-CGPE comprises three key stages: clue extraction, path exploration, and answering.
-Clue extraction serves the path exploration stage, and it cannot be experimented as an independent stage.
-For answering stage, we adopted a straightforward approach.
-Thus, we only conducted ablation studies focusing on the path exploration stage.
-The path exploration stage can be divided into three parts: character matching, large language model pruning, and termination of exploration.
-The large language model pruning includes fuzzy matching, rough matching and exact matching.
-The results of ablation studies are shown in Table 3.
-We treat direct questioning of ChatGPT as the blank control group, progressively incorporating LLM Pruning, character matching, and termination of exploration to validate the effects of each step.
+Ablation Studies
+Ablation Study on Path Exploration
+We conducted ablation studies on the path exploration stage, and the results of ablation studies are shown in Table 4.
+We treat direct answering from ChatGPT as the baseline, progressively incorporating three step matching, character matching, and termination of exploration to validate the effects of each step.
 When not using the termination of exploration, we explored each path thrice, obtaining knowledge chains with three nodes.
-In these experiments, we utilized ChatGPT for answering questions, while all other steps employed ChatGLM.
-From the table 3, it is evident that solely employing LLM pruning step yields most of the benefits of the complete process, but the invocation of the large language model remains high at 8.98 times.
+In these experiments, we utilized ChatGPT for answering questions, while all other steps employed ChatGLM3-6B.
+From the table 4, it is evident that solely employing three step matching yields most of the benefits of the complete process, but the invocation of the large language model remains high at 8.98 times.
 Introducing character matching not only improves effectiveness but also reduces the frequency of invoking the large language model.
 Choosing clues, rather than depth of exploration, as the termination criterion has no impact on effectiveness but further reduces the number of invocations of the large language model.
-As a summary, the clue-guided LLM pruning step constitutes the majority of the overall stage effectiveness, while the roles of the other two steps primarily involve reducing computational costs.
+As a summary, the three step matching constitutes the majority of the overall stage effectiveness, while the roles of the other two steps primarily involve reducing computational costs.
+Table 5: Performances with different backbone models for each stage.
+PE represents the path exploration stage, CE represents the clue extraction stage, and A represents the answering stage.
+GLM3 represents ChatGLM3-6B, CG represents ChatGPT.
 CE PE A 1-Hop Query
@@
 CG GLM CG 47.3 40.7
-G4 GLM CG 55.3 49.3
-G4 GLM G4 54.7 50.7
-GLM GLM GLM 39.6 34.2
+GPT-4 GLM3 CG 55.3 49.3
+GPT-4 GLM3 GPT-4 54.7 50.7
+GLM3 GLM3 GLM3 39.6 34.2
 CG CG CG 58.7 50.7
-G4 G4 G4 80.5 71.4
-Table 4: Performances with different backbone models for each stage.
-PE represents the path exploration stage, CE represents the clue extraction stage, and A represents the answering stage.
-GLM represents ChatGLM3-6B, CG represents ChatGPT, and G4 represents GPT-4.
-4.4.2 Performances With Different Backbone Models
+GPT-4 GPT-4 GPT-4 80.5 71.4
+Performances With Different Backbone Models
 Thanks to the plug-and-play nature of CGPE, we conducted experiments on the three stages of the framework—Clue Extraction, Path Exploration, and Answering—using different large language models with varying capabilities: ChatGLM3-6B, ChatGPT, and GPT-4.
 The purpose of these experiments is to validate the impact of models with varying capabilities on the effectiveness at each stage.
-These experiments were conducted on the Query set because, as demonstrated in Table 1, different models have a significant impact on the experimental results.
-It is conducive to showcasing the diverse influences of various models on each stage.
 The results are shown in Table 5.
@@
 However, the incremental gain is marginal when upgrading from ChatGPT to the more advanced GPT-4, as evidenced by the results on Judge sets in Table 1.
-This observation suggests that our framework may have a threshold in terms of the model’s capabilities required for the answering stage, where ChatGPT already demonstrates substantial adequacy.
+This observation suggests that ChatGPT is already capable of performing the task of “answering questions based on knowledge paths” very well.
 In the stages of clue extraction and path exploration, substituting a model with higher capabilities corresponds to superior results, with the most substantial enhancement observed in the path exploration stage.
-The path exploration stage primarily demands capabilities in information matching.
-If we fine-tune a less parameterized large language model specifically for “information matching” task Hu et al.
-(2021), it might achieve satisfactory results.
 Notably, clue extraction and answering require only one invocation of a large language model each, with the majority of computational costs concentrated in the path exploration stage.
-A less parameterized large language model for path exploration has the potential to significantly reduce the computational costs of CGPE during actual deployment.
-5 Conclusion
+Fine-tuning a less parameterized large language model specifically for “information matching” task (Hu et al.
+2021) might achieve satisfactory results while reducing the computational costs.
+Conclusion
 In this paper, we proposed a clue-guided path exploration framework named CGPE, designed to assist large language models in answering questions by exploring knowledge paths from knowledge bases.
-For a given question, CGPE initially extracts clues from the question and then explores knowledge paths guided by these clues.
-At each node along the knowledge path, we require a match with the extracted clues.
-Upon matching all clues within a knowledge path, we terminate the exploration of that path and present it to LLMs for answering the question.
-Experiments conducted on various question-answering sets indicate that CGPE outperforms previous state-of-the-art methods.
+To address the information black box issue of existing methods, we discussed the composition of necessary information and designed a path exploration mechanism for CGPE centered around information.
+All steps of the mechanism are realized through a single information matching task, reducing the requirements for the comprehensive abilities of LLMs and enhancing the ability to solve unfamiliar problems.
+Experiments on multiple datasets have demonstrated the effectiveness of our design.
+CGPE not only outperforms existing methods in terms of results, but also requires far fewer calls to LLMs than existing methods.
 Moreover, CGPE demonstrates high applicability to low-parameter LLMs compared to existing approaches, making it particularly valuable in scenarios with limited computational resources.